In Quincke's experiment the sound detected is changed from a maximum to minimum when the sliding tube is moved through a distance of 2.50cm. find the frequency of sound if the speed of sound in air is 340m s-¹

Grade:12

2 Answers

Arun
5 years ago
Dear Vikranth
 
Distance between tow maximum to a minimum is given by, λ/4 = 2.50 cm ⇒ λ = 10 cm = 10^-1 m We know, V = nx ⇒ n = V/λ = 340/10^-1 = 3400 Hz = 3.4 kHz.
 
Regards
Arun (askIITians forum expert)
KAWALDEV RAY
one year ago
Answer is correct but explanation is wrong. As when we slide tube by 2.50 cm , then path difference will be twice (2×2.50=5.0 cm).
And path difference be two consecutive maxima & minima is λ and between maxima & minima is λ/2 , hence,
λ/2 = 2 × 2.50 cm ⇒ λ = 10 cm = 10^-1 m
We know, V = nx ⇒ n = V/λ = 340/10^-1 = 3400 Hz = 3.4 kHz.